OpenStudy (jacobbirddog):

Last q for the day... medal and fan for this one!!!! Louis is planning to mow a rectangular yard that is 20 feet longer than it wide. He wants to find its width if its area is 1125 square feet. Which equation models this problem? A. 20w2 = 1125 B. w2 + 20w = 1125 C. w2 – 20w = 1125 D. w + 20w = 1125

OpenStudy (anonymous):

1125 = L * w "20 feet longer than it wide" is L = 20+W combine the two equations by replacing the L in the top eqaution with the L from the second equation done

OpenStudy (anonymous):

A = l * w l = w + 20 (20 feet longer than it is wide) A = 1125 Plug in what you know, and you'll have your answer.